OpenStudy (anonymous):

6x3/4 in radical form is cubed root of 6x to the fourth right?

OpenStudy (amistre64):

top is exponent, bottom is radical index

OpenStudy (amistre64):

is that (6x)^(3/4) or 6 x^(3/4)

OpenStudy (anonymous):

its 6x^3/4

OpenStudy (amistre64):

as is, the notation suggests that the 6 is just a multiplier to x^3/4

OpenStudy (anonymous):

\[6x^\frac{ 3 }{ 4 }\]

OpenStudy (amistre64):

\[\Large kx^{a/b}=k~\sqrt[b]{x^a}~~or~~k~(\sqrt[b]{a~})^a\]